Write the equation of the line with m=-4 and pass through (3, -1)
Answer by cira554(7)

assume that the line passes through the point(3,-1) and another point(x,y)
we know that for us to get the gradient,that is m,we get change in y divided by change in x.
so we have:y-(-1)/x-3=-4
y+1/x-3=-4
when we crossmultiply we get the equation -4(x-3)=y+1
-4x+12=y+1
y=-4x+11 which is the equation of our line